OpenStudy (anonymous):

A ball is dropped from a height of 100 ft. one second later another ball is dropped from a height of 75 ft, which one hits the ground first?

OpenStudy (anonymous):

solve \[16t^2=100\] and \[16t^2=75\] for t and compare the answers

OpenStudy (anonymous):

Okay formula is \(y = \frac{1}{2}g t^2\) For the answer 100 ft one will hit the ground first, because in one second 100 will travel 32 feet which will make the ball 68 feet away from ground.

OpenStudy (anonymous):

oh that might be a smarter way. but...

OpenStudy (anonymous):

i think the method is smarter but the conclusion may be wrong.

OpenStudy (anonymous):

oh okay at 68 feet the 100ft ball has velocity but not the 75 feet ball

OpenStudy (anonymous):

I mean one second later

OpenStudy (anonymous):

i changed my mind. now i think conclusion may be right but method is not smarter. not sure...

OpenStudy (anonymous):

\[16t^2=100\] \[t=\frac{10}{4}=2.5\] right? \[16t^2=75\] \[t=2.16\] rounded. since the first one has a second head start it landed in 1.5 seconds since the second ball was dropped

OpenStudy (anonymous):

hmm yup not a good method i let's calculate the time for each ball t for 100 feet ball 2.5 while for 75ft one \(\sqrt{4.687} +1\) the +1 is for 1 second delay and the time is 3.1 seconds 100 feet ball wins lol

OpenStudy (anonymous):

too much thinking

OpenStudy (anonymous):

can u guys give the complete solution.. ty